OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. 9x – 2y = 88 -9x + 2y = -88

OpenStudy (anonymous):

\[ a_1x+b_y=c_1\] and \[ a_2x+b_2y=c_2\] \[ \large \frac {a_1}{a_2} = \frac {b_1}{b_2}= \frac {c_1}{c_2} \implies \text{ infinite solutions }\]

OpenStudy (anonymous):

I'd like to note\[\underbrace{\begin{bmatrix}a_1&b_1\\a_2&b_2\end{bmatrix}}_A\underbrace{\begin{bmatrix}x\\y\end{bmatrix}}_B=\underbrace{\begin{bmatrix}c_1\\c_2\end{bmatrix}}_C\]\[AB=C,B=A^{-1}C\]\(A^{-1}\) can be found in multiple ways.

OpenStudy (anonymous):

For all non-infinite solution problems. :P

OpenStudy (anonymous):

Now which answer should I chose? The new medal policy isn't right :/

OpenStudy (anonymous):

ok so everyone chooses one person :D

OpenStudy (anonymous):

I choose fool and you choose bad

OpenStudy (anonymous):

(y)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i went with the first one